# Beamfocusing Optimization for Near-Field Wideband Multi-User Communications

### v4

@@ IV Low-complexity Heuristic Two-stage Approach
Although the FDA approach directly optimized the hybrid beamformer and can obtain a stationary point with the aid of the penalty method, it has the following drawbacks that hinder its practical implementation.
@@
  • Firstly, the FDA approach requires optimizing the equivalent fully-digital beamformers and the analog beamformers. In near-field communications with ELAAs, the dimension of these beamformers is typically quite large, which leads to high computational complexity for solving this optimization problem.
@@
  • Secondly, the FDA approach requires full knowledge of the CSI. Similarly, due to the large number of antenna elements in near-field communications, the complexity of the channel estimation can be extremely high.
@@
To address these issues, in this section, we propose an HTS approach, where the analog and digital beamformers are designed in two separate stages. This approach not only significantly reduces the complexity of optimization, but also simplifies channel estimation with the aid of fast near-field beam training [22]. In particular, the communication protocol including the proposed HTS approach and beam training is illustrated in Fig. 3, which will be detailed in the following.
Figure 3: Communication protocol including the proposed HTS approach and beam training.
IV-A Heuristic Analog Beamformer Design
@@
The design of the analog beamformer is based on the results of low-complexity near-field beam training [22]. The spatial wideband effect is avoided by utilizing only a small fraction of the bandwidth during the near-field beam training. We note that beam training selects the beam from a predefined codebook that has the maximum received power at the user. Given that LoS channels are much more dominant than NLoS channels, the beam obtained from beam training will be consistent with LoS channels. In other words, the approximate location of the user $(\theta_{k},r_{k})$ can be obtained by checking the location of the obtained beam in the predefined codebook. In this paper, we assume that the beam obtained by beam training perfectly matched the optimal beam to establish a performance upper-bound. Without the full CSI, the analog beamformer can be designed based on this location information related to LoS channels. In particular, the received signal of user $k$ at subcarrier $m$ through the LoS channel is given by
@@
In the following, we focus on the analog beamformer design for fully-connected architectures, which can be easily extended to cases of sub-connected architectures.
Given the user locations, the main idea of the heuristic analog beamforming design is to the analog beamformer for each RF chain to maximize the signal power for a specific user through the LoS channel [30]. Specifically, $K$ out of $N_{\mathrm{RF}}$ RF chains is used to serve the $K$ users. For fully-connected architectures, the analog beamformer at subcarrier $m$ for the $k$-th RF chain is given by
@@
where $\psi_{m,k}$ is an arbitrary phase. Unfortunately, due to the hardware limitation imposed on $\mathbf{v}_{m,k}$, it is generally impossible to find $\mathbf{a}_{k,l}$ and $t_{k,l},\forall l,$ such that $\mathbf{v}_{m,k}=\mathbf{v}_{m,k}^{\mathrm{opt}},\forall m$. As a remedy, in the following, we propose a high-quality piecewise-near-field approximation of the near-field array response vector. Then, a computationally efficient design of the analog beamformer is proposed based on the piecewise-near-field approximation.
@@
IV-A1 Piecewise-Near-Field Approximation
This approximation is based on the fact that the entire antenna array is divided into $N_{\mathrm{T}}$ sub-arrays by TTDs. Therefore, we divide the array response vector $\mathbf{b}_{N}(f_{m},\theta,\nu)$ into $N_{\mathrm{T}}$ sub-vectors corresponding to $N_{\mathrm{T}}$ TTDs as follows:
@@ IV-A1 Piecewise-Near-Field Approximation
The relationship between the worst-case accuracy of the piecewise-far-field approximation and the number of TTDs is given in the following proposition.
Figure 5: Illustration of function $\delta\left(\frac{N_{\mathrm{T}}}{N},\frac{B}{f_{c}}\right)$.
@@
According to Proposition 1, the accuracy of the piecewise-near-field approximation in the worst-case scenario depends on two key ratios: $N_{\mathrm{T}}/N$ and $B/f_{c}$. The behavior of the function $\delta\left(\frac{N_{\mathrm{T}}}{N},\frac{B}{f_{c}}\right)$ in relation to these ratios is depicted in Fig. 5. It is evident that an increase in bandwidth $B$ necessitates a corresponding increase in the number of TTDs to maintain a specified accuracy threshold $\Delta$. For instance, consider a scenario with $\Delta=0.8$, $N=512$, and $f_{c}=100$ GHz. In this context, as per the criterion given in (61), a bandwidth of $B=10$ GHz requires $N_{\mathrm{T}}\geq 36$ TTDs. Conversely, for a larger bandwidth of $B=20$ GHz, the requirement escalates to $N_{\mathrm{T}}\geq 105$ TTDs. It is important to note that the above analysis pertains to the worst-case performance. When considering the average performance, the number of TTDs required can be significantly lower, which will be demonstrated by the numerical results presented in Section V.
1:  Stage I: Heuristic analog beamformer design
2:  estimate the user location information $\theta_{k}$ and $\nu_{k},\forall k,$ through low-complexity near-field beam training.
3:  for  $k\in\{1,\dots,N_{\mathrm{RF}}\}$  do
4:     If $n\leq K$, design $\mathbf{a}_{n,l},\forall l,$ according to (64), and design $t_{n,l},\forall l,$ as the solution to (IV-A2). Otherwise, design them based on the location of a randomly selected user.
5:  end for
6:  Stage II: Digital beamformer optimization
7:  estimate the equivalently channel $\tilde{\mathbf{h}}_{m,k},\forall m,k$.
8:  optimize the digital beamformer $\mathbf{D}_{m},\forall m,$ by solving (IV-B).
IV-A2 Analog Beamformer Design
@@
$\bar{G}(\theta_{k},r_{k})=\frac{1}{M}\sum_{m\in\mathcal{M}}\big{|}\hat{\mathbf{b}}^{T}(f_{m},\theta_{k},r_{k})\mathbf{v}_{m,k}\big{|}$
$=\frac{1}{M}\sum_{m\in\mathcal{M}}\left|\sum_{l=1}^{N_{\mathrm{T}}}\hat{\bm{\varphi}}_{l}^{T}(\theta_{k},r_{k})\mathbf{a}_{k,l}e^{-j2\pi f_{m}\left(\frac{\nu_{l}(\theta_{k},r_{k})-r_{k}}{c}+t_{k,l}\right)}\right|.$ (62)
@@
The above optimization problem is essentially a phase alignment problem. According to (IV-A2), to maximize $\bar{G}(\theta_{k},r_{k})$, the PS-based analog beamformer $\mathbf{a}_{k,l}$ can be designed to match the phase of the frequency-independent vector $\hat{\bm{\varphi}}_{l}^{T}(\theta_{k},r_{k})$, resulting the following solution:
@@
Given the above solution, problem (IV-A2) reduces to
$\max_{t_{k,l}}\quad$ $\sum_{m\in\mathcal{M}}\left|\sum_{l=1}^{N_{\mathrm{T}}}e^{-j2\pi f_{m}\left(\frac{\nu_{l}(\theta_{k},r_{k})-r_{k}}{c}+t_{k,l}\right)}\right|$ (65a)
@@
Although the optimization variables are highly coupled in the objective function, this problem can be solved through an iterative one-dimensional search. More specifically, in each iteration, each time delay $t_{k,l}$ is updated by one-dimensional search as given in (30) and (45) by fixing the other time delays. The complexity of the iterative one-dimensional search is on the order of $\mathcal{O}(I_{\text{iter}}MN_{\mathrm{T}})$.
IV-B Digital Beamformer Optimization
@@
where $\tilde{\mathbf{h}}_{m,k}=\mathbf{T}_{m}^{H}\mathbf{A}^{H}\mathbf{h}_{m,k}$ represents the equivalent channel. This equivalent channel, which has a low dimensionality of $N_{\mathrm{RF}}$, can be efficiently estimated using standard channel estimation methods with low complexity. The digital beamformer can then be optimized to maximize the spectral efficiency of the equivalent channel, resulting in the following optimization problem:
$\max_{\mathbf{D}_{m}}\quad$ $\sum_{m\in\mathcal{M}}\sum_{k\in\mathcal{K}}\breve{R}_{m,k}(\mathbf{D}_{m})$ (67a)
@@
This problem is a classical sum-rate maximization, which can be effectively solved by the existing method, such as fractional programming [29], successive convex approximation [31], and weighted minimum mean-squared error [32] methods. Thus, we omit the details here. Due to the low dimensionality of $\mathbf{D}_{m}$, the corresponding algorithm has low complexity. Based on the above discussion, the overall HTS approach is summarized in Algorithm 4.
For sub-connected architectures, the HTS approach follows a similar procedure as in Algorithm 4. The difference is that the design of the analog beamformer needs to be based on a piecewise-near-field approximation of the sub-arrays connected to each RF chain.

### v6

@@ IV Heuristic Two-stage Approach
Although the FDA approach directly optimizes the hybrid beamformer and can obtain a stationary point with the aid of the penalty method, it has the following drawbacks.
@@
  • First, the FDA approach requires the joint optimization of the equivalent fully-digital beamformers and the analog beamformers. In near-field communications with ELAAs, the dimension of these beamformers is typically quite large, which leads to relatively high computational complexity for solving this optimization problem.
@@
  • Second, the FDA approach requires full knowledge of the CSI. Similarly, due to the large number of antenna elements in near-field communications, the complexity of the channel estimation can be extremely high.
@@
To address these issues, in this section, we propose an HTS approach, where the analog and digital beamformers are designed in two separate stages. At the cost of slight performance loss, this approach not only significantly reduces the complexity of optimization, but also simplifies channel estimation with the aid of fast beam training or beam tracking [23, 37]. In particular, the communication protocol including the proposed HTS approach and beam training is illustrated in Fig. 3. The advantages of this protocol in reducing the computational complexity and signaling overhead of channel estimation have been widely demonstrated in the literature [5, 38]. Thus, we only focus on the beamforming design in the following.
IV-A Heuristic Analog Beamforming Optimization
@@
The design of the analog beamformer is based on the results of low-complexity near-field beam training or tracking [23, 37]. The spatial wideband effect can be avoided by utilizing only a small fraction of the bandwidth during beam training or tracking. We note that beam training selects the beam from a predefined codebook that has the maximum received power at the user. Given that LoS channels are much more dominant than NLoS channels, the beam obtained from beam training will be consistent with LoS channels. In other words, the approximate location of the user $(\theta_{k},r_{k})$ can be obtained by checking the location of the obtained beam in the predefined codebook. Furthermore, beam tracking is an ongoing process that adjusts the beam in response to changes in the environment or the movement of the user, thus avoiding the searching process in beam training. In this paper, we assume that the beam obtained by beam training or tracking perfectly matched the optimal beam to establish a performance upper bound. Without the full CSI, the analog beamformer can be designed based on this location information related to LoS channels. In particular, the received signal of user $k$ on subcarrier $m$ through the LoS channel is given by
@@
In the following, we focus on the analog beamforming design for fully-connected architectures, which can be easily extended to cases of sub-connected architectures.
Figure 3: Communication protocol including the proposed HTS approach and beam training/tracking.
Given the user locations, the main idea of the heuristic analog beamforming design is to the analog beamformer for each RF chain to maximize the signal power for a specific user through the LoS channel [39]. Specifically, $K$ out of $N_{\mathrm{RF}}$ RF chains is used to serve the $K$ users. For fully-connected architectures, the analog beamformer on subcarrier $m$ for the $k$-th RF chain is given by
@@
where $\psi_{m,k}$ is an arbitrary phase. Unfortunately, due to the hardware limitation imposed on $\mathbf{v}_{m,k}$, it is generally impossible to find $\mathbf{a}_{k,l}$ and $t_{k,l},\forall l,$ such that $\mathbf{v}_{m,k}=\mathbf{v}_{m,k}^{\mathrm{opt}},\forall m$. As a remedy, in the following, we propose two low-complexity analog beamforming designs based on the approximated and exact LoS channels, respectively.
@@
IV-A1 Piecewise-Near-Field (PNF)-based Design
This design employs a PNF approximation of the LoS channels. In practice, the PNF approximation is based on the fact that the entire antenna array is divided into $N_{\mathrm{T}}$ sub-arrays by TTDs. Therefore, we divide the array response vector $\mathbf{b}_{N}(f_{m},\theta,\nu)$ into $N_{\mathrm{T}}$ sub-vectors corresponding to $N_{\mathrm{T}}$ TTDs as follows:
@@ IV-A1 Piecewise-Near-Field (PNF)-based Design
The relationship between the worst-case accuracy of the PNF approximation and the number of TTDs is given in the following proposition.
Figure 5: Illustration of function $\delta\big{(}\frac{N_{\mathrm{T}}}{N},\frac{B}{f_{c}}\big{)}$.
@@
According to Proposition 1, the accuracy of the PNF approximation in the worst-case scenario depends on two key ratios: $N_{\mathrm{T}}/N$ and $B/f_{c}$. The behavior of the function $\delta\big{(}\frac{N_{\mathrm{T}}}{N},\frac{B}{f_{c}}\big{)}$ in relation to these ratios is depicted in Fig. 5. It is evident that an increase in bandwidth $B$ necessitates a corresponding increase in the number of TTDs to maintain a specified accuracy threshold $\Delta$. For instance, consider a scenario with $\Delta=0.8$, $N=512$, and $f_{c}=100$ GHz. In this context, as per the criterion given in (57), a bandwidth of $B=10$ GHz requires $N_{\mathrm{T}}\geq 36$ TTDs. Conversely, for a larger bandwidth of $B=20$ GHz, the requirement escalates to $N_{\mathrm{T}}\geq 105$ TTDs. It is important to note that the above analysis pertains to the worst-case performance. When considering the average performance, the number of TTDs required can be significantly lower, which will be demonstrated by the numerical results presented in Section V.
@@
$\bar{G}(\theta_{k},r_{k})=\frac{1}{M}\sum_{m=1}^{M}\big{|}\hat{\mathbf{b}}^{T}(f_{m},\theta_{k},r_{k})\mathbf{v}_{m,k}\big{|}$
$=\frac{1}{M}\sum_{m=1}^{M}\left|\sum_{l=1}^{N_{\mathrm{T}}}\hat{\bm{\varphi}}_{l}^{T}(\theta_{k},r_{k})\mathbf{a}_{k,l}e^{-j2\pi f_{m}\left(\frac{\nu_{l}(\theta_{k},r_{k})-r_{k}}{c}+t_{k,l}\right)}\right|.$ (58)
@@
The above optimization problem is essentially a phase alignment problem. According to (IV-A1), to maximize $\bar{G}(\theta_{k},r_{k})$, the PS-based analog beamformer $\mathbf{a}_{k,l}$ can be designed to match the phase of the frequency-independent vector $\hat{\bm{\varphi}}_{l}^{T}(\theta_{k},r_{k})$, resulting the following solution:
@@
Given the above solution, problem (59) reduces to
$\max_{t_{k,l}}\quad$ $\sum_{m=1}^{M}\left|\sum_{l=1}^{N_{\mathrm{T}}}e^{-j2\pi f_{m}\left(\frac{\nu_{l}(\theta_{k},r_{k})-r_{k}}{c}+t_{k,l}\right)}\right|$ (61a)
@@
For the solution to this optimization problem, we have the following proposition.
Proposition 2.
If $t_{\max}\geq\frac{N(N_{\mathrm{T}}-1)d}{N_{\mathrm{T}}c}$, the optimal solution to problem (61) is given by
$t_{k,l}=\frac{\nu_{\max,k}-\nu_{l}(\theta_{k},r_{k})}{c},$ (62)
where $\nu_{\max,k}=\max\{\nu_{1}(\theta_{k},r_{k}),\dots,\nu_{N_{\mathrm{T}}}(\theta_{k},r_{k})\}$.
Proof:
Please refer to Appendix B. ∎
@@
In Proposition 2, we prove that the optimal TTD coefficients to problem (61) have the closed-form solution when $t_{\max}$ is sufficiently large. When $t_{\max}$ is small, problem (61) can be solved through an iterative one-dimensional search. More specifically, in each iteration, each time delay $t_{k,l}$ is updated by one-dimensional search as given in (30) and (44) by fixing the other time delays. The complexity of the one-dimensional search in each iteration is $\mathcal{O}(QMN_{\mathrm{T}})$.
IV-A2 Robust Design
Although the idea of the PNF-based design can facilitate the closed-form solution of the analog beamformers, it might result in substantial performance degradation under certain conditions. Specifically, this design relies extensively on the accuracy of the PNF approximation of the channel responses. However, this accuracy can diminish significantly when the number of TTDs $N_{\mathrm{T}}$ is small, or when the system operates over an ultra-large bandwidth $B$, as illustrated in Fig. 5. To address this issue, in this section, we further propose a robust design that jointly optimizes the PS-based and TTD-based analog beamformers based on the exact LoS channel. The exact average array gain achieved by the overall analog beamformer for user $k$ is given by
$\tilde{G}(\theta_{k},r_{k})=$ $\frac{1}{M}\sum_{m=1}^{M}\left|\mathbf{b}^{T}(f_{m},\theta_{k},r_{k})\mathbf{v}_{m,k}\right|$
$=$ $\frac{1}{M}\sum_{m=1}^{M}\left|\sum_{l=1}^{N_{\mathrm{T}}}\bm{\varphi}_{m,l}^{T}(\theta_{k},r_{k})\mathbf{a}_{k,l}e^{-j2\pi f_{m}t_{k,l}}\right|,$ (63)
where $\bm{\varphi}_{m,l}$ is defined in (48). In the sequel, we jointly optimize the PS-based and TTD-based analog beamformers to maximize the exact average array gain, resulting in the following optimization problem:
$\max_{\mathbf{a}_{k,l},t_{k,l}}\quad$ $\sum_{m=1}^{M}\left|\sum_{l=1}^{N_{\mathrm{T}}}\bm{\varphi}_{m,l}^{T}(\theta_{k},r_{k})\mathbf{a}_{k,l}e^{-j2\pi f_{m}t_{k,l}}\right|$ (64a)
$\mathrm{s.t.}\quad$ $|\mathbf{a}_{k,l}(i)|=1,\forall l,i,$ (64b)
$t_{k,l}\in[0,t_{\max}],\forall l.$ (64c)
This problem can also be solved using BCD given as follows.
When $t_{k,l}$ is fixed, problem (64) can be reformulated as
$\max_{\bar{\mathbf{a}}_{k}}\quad$ $\sum_{m=1}^{M}\left|\bm{\eta}_{m,k}^{H}\bar{\mathbf{a}}_{k}\right|$ (65a)
$\mathrm{s.t.}\quad$ $|\bar{\mathbf{a}}_{k}(n)|=1,\forall n,$ (65b)
where $\bar{\mathbf{a}}_{k}=[\mathbf{a}_{k,1}^{T},\dots,\mathbf{a}_{k,N_{\mathrm{T}}}^{T}]^{T}$ and
$\bm{\eta}_{m,k}=\begin{bmatrix}\bm{\varphi}_{m,1}^{*}(\theta_{k},r_{k})e^{j2\pi f_{m}t_{k,1}}\\ \vdots\\ \bm{\varphi}_{m,N_{\mathrm{T}}}^{*}(\theta_{k},r_{k})e^{j2\pi f_{m}t_{k,N_{\mathrm{T}}}}\end{bmatrix}.$ (66)
This optimization problem is non-convex. We propose a low-complexity iterative algorithm to solve it based on the majorization-minimization (MM) method [40]. In particular, we first construct the following surrogate function of the objective function:
$f\left(\bar{\mathbf{a}}_{k}|\bar{\mathbf{a}}_{k}^{(t)}\right)\triangleq\sum_{m=1}^{M}\Re\left\{\frac{\left(\bar{\mathbf{a}}_{k}^{(t)}\right)^{H}\bm{\eta}_{m,k}\bm{\eta}_{m,k}^{H}\bar{\mathbf{a}}_{k}}{\left|\bm{\eta}_{m,k}^{H}\bar{\mathbf{a}}_{k}^{(t)}\right|}\right\},$ (67)
where $\bar{\mathbf{a}}_{k}^{(t)}$ is the solution of $\bar{\mathbf{a}}_{k}$ obtained in the $t$-th iteration of the MM algorithm. According to the Cauchy-Schwarz inequality, we have
$f\left(\bar{\mathbf{a}}_{k}|\bar{\mathbf{a}}_{k}^{(t)}\right)\leq\sum_{m=1}^{M}\left|\bm{\eta}_{m,k}^{H}\bar{\mathbf{a}}_{k}\right|,\vspace{-0.1cm}$ (68)
where the equality is achieved at $\bar{\mathbf{a}}_{k}^{(t)}=\bar{\mathbf{a}}_{k}$. With the above surrogate function $f\left(\bar{\mathbf{a}}_{k}|\bar{\mathbf{a}}_{k}^{(t)}\right)$, problem (65) can be solved by the following process [40]:
$\bar{\mathbf{a}}_{k}^{(t+1)}=\operatorname*{arg\,max}_{|\bar{\mathbf{a}}_{k}(n)|=1,\forall n}f\left(\bar{\mathbf{a}}_{k}|\bar{\mathbf{a}}_{k}^{(t)}\right)=e^{j\angle\left(\mathbf{q}_{k}^{(t)}\right)},$ (69)
where
$\mathbf{q}_{k}^{(t)}=\sum_{m=1}^{M}\frac{\bm{\eta}_{m,k}\bm{\eta}_{m,k}^{H}\bar{\mathbf{a}}_{k}^{(t)}}{\left|\bm{\eta}_{m,k}^{H}\bar{\mathbf{a}}_{k}^{(t)}\right|}.$ (70)
After obtaining $\bar{\mathbf{a}}_{k}$ through the iterative MM algorithm, the PS-based analog beamformer can be designed as $\mathbf{a}_{k,l}=\bar{\mathbf{a}}_{k}\big{(}(l-1)\frac{N}{N_{\mathrm{T}}}+1:l\frac{N}{N_{\mathrm{T}}}\big{)},\forall l$.
When $\mathbf{a}_{k,l}$ is fixed, problem (64) reduces to
$\max_{t_{k,l}}\quad$ $\sum_{m=1}^{M}\left|\sum_{l=1}^{N_{\mathrm{T}}}\gamma_{m,k,l}e^{-j2\pi f_{m}t_{k,l}}\right|$ (71a)
$\mathrm{s.t.}\quad$ $t_{k,l}\in[0,t_{\max}],\forall l,$ (71b)
where $\gamma_{m,k,l}=\bm{\varphi}_{m,l}^{T}(\theta_{k},r_{k})\mathbf{a}_{k,l}$. Similar to problem (61), this problem can also be solved through the iterative one-dimensional search. Then, the solution to problem (64) can be obtained by updating $\mathbf{a}_{k,l}$ and $t_{k,l}$ alternately based on the above methods. The complexity for updating $\mathbf{a}_{k,l},\forall l,$ in each iteration of the MM algorithm is $\mathcal{O}(2MN^{2}+MN)$ and for updating $t_{k,l},\forall l$, through the one-dimensional search is $\mathcal{O}(QMN_{\mathrm{T}})$. Compared to the PNF-based design, the robust design maximizes the array gain based on the exact channel model, resulting in a better performance but at the cost of incurring higher complexity.
IV-B Digital Beamforming Optimization
@@
where $\tilde{\mathbf{h}}_{m,k}=\mathbf{T}_{m}^{H}\mathbf{A}^{H}\mathbf{h}_{m,k}$ represents the equivalent channel. This equivalent channel, which has a low dimensionality of $N_{\mathrm{RF}}$, can be efficiently estimated using standard channel estimation methods with low complexity. The digital beamformer can then be optimized to maximize the spectral efficiency of the equivalent channel, resulting in the following problem:
$\max_{\mathbf{D}_{m}}\quad$ $\sum_{m=1}^{M}\sum_{k=1}^{K}\breve{R}_{m,k}(\mathbf{D}_{m})$ (73a)
@@
This problem is a classical sum-rate maximization, which can be effectively solved by the existing method, such as fractional programming [34], MM [40], and weighted minimum mean-squared error [41] methods. Thus, we omit the details here. Due to the low dimensionality of $\mathbf{D}_{m}$, the corresponding algorithm has low complexity. Based on the above discussion, the overall HTS approach is summarized in Algorithm 4.
For sub-connected architectures, the HTS approach follows a similar procedure as in Algorithm 4. The difference is that the design of the analog beamformer needs to be based on the array response of the sub-arrays connected to each RF chain.
1:  Stage I: Heuristic analog beamforming design
2:  estimate the user location information $\theta_{k}$ and $r_{k},\forall k$.
3:  for  $n\in\{1,\dots,N_{\mathrm{RF}}\}$  do
4:     If $n\leq K$, design $\mathbf{a}_{n,l}$ and $t_{n,l},\forall l$ based on the location of user $n$, i.e., $\theta_{n}$ and $r_{n}$, using the PNF-based method or the robust method. Otherwise, design them based on the location of a randomly selected user.
5:  end for
6:  Stage II: Digital beamforming optimization
7:  estimate the equivalent channel $\tilde{\mathbf{h}}_{m,k},\forall m,k$.
8:  optimize the digital beamformer $\mathbf{D}_{m},\forall m,$ by solving (73).